You are given a cylinder, a ring and a sphere with the same mass and radii respectively. Which has the least moment of inertia?
Answer #1
  1. moment of inertia- 1.Ring- I=Mr2​​​​​
  2. Hollow ​​cylinder- I=Mr2
  3. solid cylinder- I=1/2Mr2
  4. hollow  sphere- I=2/3Mr2
  5. solid sphere- I=2/5Mr2 Then solid spare has the least moment of inertia